In 2012, azania had nominal gdp of $6.6 million. Inflation rate with 2011 as the base year was 105. in 2013 the country's nominal gdp increased to $6.9 million with inflation rate of 120 on the basis of the same 2011 base year. what was the growth of real gdp between 2012 and 2013?
1
Expert's answer
2014-11-25T10:13:46-0500
GDPn(2012) = $6.6m, CPI(2012) = 105, GDPn(2013) = $6.9m, CPI(2013) = 120
Real GDP = Nominal GDP/CPI*100
Real GDP(2012) = 6.6/105*100 = $6.29 million
Real GDP(2012) = 6.9/120*100 = $5.75 million
Real GDP growth rate = (5.75 - 6.29)/6.29*100% = -8.6%